Twice the sum of a number and -6 equals -6 times the sum of the number and 4. Let n stand for the number. Solve for n
Answer by RayVelbis(4)

2n-6=-6+n(4)
2n-6=-6+4n
-6=-6+2n
NO SOLUTION